OpenStudy (anonymous):

how to solve 5/7x - 8/9

OpenStudy (anonymous):

This is merely an expression and thus has no solution. What does the problem say?

OpenStudy (phi):

you can solve equations... equations have an = sign in them

OpenStudy (anonymous):

it says to simplify

OpenStudy (phi):

is the problem \[ \frac{5}{7x} - \frac{8}{9} \]?

OpenStudy (anonymous):

The expression given is already simplified as far as I would simplify it

OpenStudy (anonymous):

5 / 7x- 8/ 9 =0 5*9=8*7x 8*7x=5*9 56x=5*9 56x=45 x= 45 / 56

OpenStudy (anonymous):

The word " solve" should be an indicator that the poster forgot to equate to 0